Formulate the demand equations and estimate Qd for P=33 by using the following data:   


Price level Quantity Demand

38 200

36 500

34 800

32 900

30 1000

28 1400


1
Expert's answer
2021-04-27T07:43:41-0400

The formula for demand equation:

"y-y1=m(x-x1)"

y is the dependent variable which is price

x is the independent variable which is the quantity demanded

m is the slope

"m=\frac{y2-y1}{x2-x1}"

Let us take two points from the given data:

"y1=P1=36"

"x1=Q1=500"

"y2=P2=32"

"x2=Q2=900"

"m=\frac{P2-P1}{Q2-Q1}"

"m=\frac{32-36}{900-500}"

"m=\frac{-4}{400}=-0.01"

A negative slope means the negative relationship between quantity demand and price


"P-P1=m(Q-Q1)"

"P-36=-0.01(Q-500)"

"P-36=-0.01Q+5"

 "P=41-0.01Q" is the damand equition


When P=33,

"33=41-0.01Q"

"Q=\frac{33-41}{-0.01}"

"Q=800"

 

When price is 33, quantity demanded is 800 units.